OpenStudy (anonymous):

Solve: 2(x + 1) = −8

OpenStudy (anonymous):

distribute first!

OpenStudy (homeworksucks):

so you distribute first 2x+2=-8 combine like terms 2x=-10 x=-5

OpenStudy (anonymous):

2(x+1) = -8 (x+1) = -8/2 (x+1) = -4 x = -4-1 x = -5

OpenStudy (anonymous):

Please no giving answers away guys!!!

OpenStudy (anonymous):

You can be banned for that :/

OpenStudy (anonymous):

thanks

OpenStudy (aravindg):

well @RyanL. has a point there .It would be a beeter idea if you help the asker get to the solution instead of directly posting the answers .read the coc http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

What I do is always leave the last step for the person asking the question.

OpenStudy (anonymous):

Usually it's one of the easier steps :D

OpenStudy (aravindg):

lols

OpenStudy (anonymous):

sorry to ask inside ur problem roberto.. but does anybody know how to do the echelon method to solve the system of equations?

OpenStudy (homeworksucks):

oooh sorry guys didn't know... I'll try not to next time